# MoLA: Motion Generation and Editing with Latent Diffusion Enhanced by Adversarial Training

### v2

@@ 3.2.1. Learning a rich motion latent representation with VAE-GAN
where $\lambda_{\text{reg}}$ is a reciprocal of $\sigma^{2}$ (treated as a hyperparameter in this work). The encoder $q_{\bm{\eta}}$ can be trained to produce a low-dimensional latent representation, and the decoder $g_{\bm{\psi}}$ can also be trained to accurately reconstruct the input motion sequences from the latent representations.
@@
We formulate the overall loss for the VAE as a combined objective of the negative ELBO and adversarial loss, as
(3) $\min_{\phi,\psi}\mathbb{E}_{p(x)}\left[\mathcal{J}_{\text{VAE}}(\bm{\psi},\bm{\eta};\bm{x})+\lambda_{\text{adv}}\mathcal{J}_{\text{GAN}}(\bm{\psi},\bm{\eta};\bm{\phi},\bm{x})\right],$
@@
We train both the VAE and the discriminator with Equations (2) and (3) in an alternating way.
@@ 3.2.2. From GAN- to SAN-based discriminator
We apply the SAN framework (Takida et al., 2024) to further enhance the motion VAE based on a prior report showing that a model with a SAN-based discriminator has achieved SOTA performance in image generation. Takida et al. (Takida et al., 2024) incorporated a sliced optimal transport perspective into a general GAN. Following this work, we first decompose the discriminator $f_{\bm{\phi}}$ into the last linear layer and the remaining neural part, denoted as $\bm{w}\in\mathbb{R}^{d_{w}}$ and $h_{\bm{\varphi}}:\mathbb{R}^{N\times L}\to\mathbb{R}^{d_{w}}$ with $d_{w}\in\mathbb{N}$, respectively. This decomposition provides an interpretation of the discriminator: the neural function $h_{\bm{\varphi}}$ maps motion sequences to non-linear features, and then the linear layer $\bm{w}$ projects them into scalars. As a preliminary step for applying SAN to the adversarial training in Section 3.2.1, we normalize $w$ using its norm, resulting in $\bm{\omega}=\bm{w}/\|\bm{w}\|_{2}$, which indicates the direction of the projecting. Now, the discriminator is represented in the form of an inner product as $f_{\bm{\phi}}(\bm{x})=\bm{\omega}^{\top}h_{\varphi}(\bm{x})$, and its parameter is $\bm{\phi}=\{\bm{\varphi},\bm{\omega}\}$.
$\mathcal{L}_{\text{SAN}}(\bm{\phi};\bm{\psi},\bm{\eta},\bm{x})$ $=\mathcal{L}_{\text{GAN}}(\{\bm{\varphi},\bm{\omega}^{-}\};\bm{\psi},\bm{\eta},\bm{x})$
(5) $+\bm{\omega}^{\top}\left(h_{\bm{\varphi}^{-}}(\bm{x})-\mathbb{E}_{q_{\bm{\eta}}(\bm{z}|\bm{x})}[h_{\bm{\varphi}^{-}}(g_{\bm{\psi}}(\bm{z}))]\right),$
@@ 3.3. Stage 2: Motion Latent Diffusion
In this section, we train a text-conditioned diffusion model on the low-dimensional motion latent space obtained by the autoencoder learned in the stage 1 (Section 3.2). Using the trained model, we perform motion generation conditioned on text. First, we define a time-dependent sequence $\bm{z}=\bm{z}_{0},\bm{z}_{1},\dots,\bm{z}_{t},\dots,\bm{z}_{T}\in\mathbb{R}^{d_{z}}$ starting from the VAE encoder output $\bm{z}\sim q_{\bm{\eta}}(\bm{z}|\bm{x})$, which is derived from the following Markov diffusion process in the latent space:
@@
where $T$ 0 and the constant $\alpha_{t}\in(0,1)$ is a pre-defined noise-scheduling parameter that determines the forward process. The forward process allows for the sampling of $\bm{z}_{t}$ at an arbitrary time step $t$ in a closed form: $\bm{z}_{t}=\sqrt{\bar{\alpha}_{t}}\bm{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$, where $\bar{\alpha}_{t}\coloneq\prod_{s=1}^{t}\alpha_{s}$ and $\epsilon\sim\mathcal{N}(\bm{0},\bm{I})$.
The diffusion model is trained to reverse the diffusion process described in Equation (6), i.e., to perform denoising. Specifically, the model learns to estimate the noise component $\epsilon$ in the noisy sample $\bm{z}_{t}$ as in the following objective:
(7) $\mathcal{J}_{\text{LDM}}(\theta)=\mathbb{E}_{\bm{z}_{0},\epsilon,t}\left[\|\epsilon-\epsilon_{\theta}(\bm{z}_{t},t)\|_{2}^{2}\right],$
where $\theta$ is the model parameter to be trained, and the time step $t$ is uniformly sampled from $\{1,\dots,T\}$.
As our goal is text-to-motion generation, our interest is in the conditional distribution $p(\bm{z}|\bm{c})$ conditioned on the text prompt $\bm{c}$. Here, similar to many text-conditioned latent diffusion models (Rombach et al., 2022; Chen et al., 2023), we train the conditional model $\epsilon_{\theta}(\bm{z}_{t},t,\tau(\bm{c}))$ conditioned on the output of a text encoder $\tau(\bm{c})$, using the following objective function:
@@
where $\bm{z}_{0}$ and $\bm{c}$ are drawn from the joint empirical distribution. In addition, following many other works, we adopt classifier-free guidance (Ho and Salimans, 2022) and train the model unconditionally i.e., without a text prompt, with a certain probability during training.
During the inference (text-to-motion generation), the trained diffusion model $\epsilon_{\theta}(\bm{z}_{t},t,\tau(\bm{c}))$ is employed to generate $\bm{z}_{0}$ through a denoising process conditioned on the text prompt $\bm{c}$. We adopt the sampling scheme of DDIM (Song et al., 2021a), in which each sampling step is defined as:
@@ 3.4. Controllable Motion Generation on Latent Diffusion Sampling
as derived from Bayes’ rule. The conditional generation based on the aforementioned property can be regarded as a sequential procedure implemented as follows. First, from $\bm{z}_{t}$, a denoised sample $\bm{z}_{t-1}$ is obtained by the sampling step in Equation (9), without considering the given condition $\bm{y}$. Subsequently, $\bm{z}_{t-1}$ is further updated using the gradient with respect to $\bm{z}_{t}$ for the log-likelihood term.
The log-likelihood term is defined for noisy sample $\bm{z}_{t}$. In the classifier-guided diffusion (Song et al., 2021b), time-dependent classifiers for $\bm{z}_{t}$ have to be trained, which requires additional training. In contrast, in training-free methods such as DPS, FreeDoM, and MPGD, this term is approximated with the current clean data estimate $\bm{z}_{0|t}$ and a loss function $L(\bm{x};\bm{y})$ defined for clean data. This loss function can be flexibly set depending on the task. For example, in inverse problems of the form $\bm{y}=\mathcal{A}(\bm{x})$, where $\mathcal{A}$ is a differentiable function with respect to $\bm{x}$, the loss function can be set as $\|\bm{y}-\mathcal{A}(\bm{x}_{0|t})\|_{2}^{2}$, where $\bm{x}_{0|t}=g_{\bm{\psi}}(\bm{z}_{0|t})$ is a clean data estimate in the original data domain and obtained through the VAE decoder. In FreeDoM’s FaceID-guided generation, the $\ell_{2}$ distance of features obtained by inputting facial images into a facial recognition model is used as the loss.
Here, we adopt MPGD (He et al., 2024), a fast yet high-quality guidance method applicable to latent diffusion models. Following the denoising step, it updates the denoised sample for the loss function $L(\bm{x};\bm{y})$ as follows:
(11) $\bm{z}_{t-1}\leftarrow\bm{z}_{t-1}-\rho_{t}\sqrt{\bar{\alpha}_{t-1}}\nabla_{\bm{z}_{0|t}}L(g_{\bm{\psi}}(\bm{z}_{0|t});\bm{y}),$
@@
More specifically, in editing motion tasks we generate motions to match given specific poses or a trajectory control signal. To deal with various motion editing tasks in this framework, the loss functions are designed as follows:
(12) $L(g_{\bm{\psi}}(\bm{z}_{0|t});\bm{y})=\Sigma_{n}\Sigma_{l}m_{nl}\|R(g_{\bm{\psi}}(\bm{z}_{0|t}))_{nl}-y_{nl}\|_{2},$
@@
where $n$ and $l$ are indices of joint and frame, respectively, and $m_{nl}$ is a binary value indicating whether the control position $y_{nl}$ contains a valid value at frame $l$ for joint $n$, and $R(\cdot)$ is a function to convert the motion features including the joint’s local positions to global absolute locations. We set $L$ for the guided generation to measure the distance between desired constraints $\bm{y}$ and the joint locations of the generated motion. Target locations as constraint $\bm{y}$ can be specified for any subset of joints in any subset of motion frames. [^1] As examples of motion editing, if $\bm{y}$ is given as the start-end positions, we can handle the motion in-betweening. If $\bm{y}$ is given as the lower body positions, we can edit the upper body corresponding to the lower one. If $\bm{y}$ is given as the pelvis trajectory, it corresponds to the path-following task. We leave the task details to Section 4.5. Note that the guided generation framework in Equations (11) and (12) has the potential to generate motion while dealing with a variety of time and spatial constraints not limited to these three task examples. Editing a generated motion to match specific poses or to follow a specific trajectory can be achieved by minimizing $L$ with the update rule in Equation (11).
4. Experiments
4.1. Datasets
We conduct empirical evaluations on two widely used datasets for text-to-motion generation: HumanML3D (Guo et al., 2022) and KIT Motion-Language (KIT-ML) (Plappert et al., 2016). HumanML3D contains 14,616 human motions from the AMASS (Mahmood et al., 2019) and HumanAct12 (Guo et al., 2020) datasets and 44,970 text descriptions. KIT-ML contains 3,911 human motion sequences from the KIT (Mandery et al., 2015) and CMU (CMU, 2003) datasets and 6,278 textual descriptions. Both datasets are split into training, validation, and test sets with proportions of 80%, 5% and 15%, respectively.
Dataset Stage 1 model rFID $\downarrow$
HumanML3D VAE 0.206
VAE + GAN 0.179
VAE + SAN 0.141
KIT-ML VAE 0.311
VAE + GAN 0.293
VAE + SAN 0.280
Table 2. Analysis of motion reconstruction performance (training stage 1 with VAE) on HumanML3D and KIT-ML
4.2. Implementation Details
In our experiment, we train two different types of models, as follows. (i) We train the motion VAE and discriminator $f_{\bm{\psi}}$ described in Section 3.2 and use these stage 1 models to train the stage 2 model $\epsilon_{\theta}$. (ii) We train the motion VQ-VAE in T2M-GPT (Zhang et al., 2023b) and the discriminator, and train the transformer model in the latter stage to generate code indices conditioned on the text description. Case (ii) is intended to verify whether adversarial training is also effective in a VQ-based text-to-motion framework.
In case (i), the motion VAE architecture is based on the VAE part in MLD (Chen et al., 2023). The encoder $q_{\bm{\eta}}$ and decoder $g_{\bm{\psi}}$ of the VAE model all consist of nine layers and four heads with skip connection. For adversarial training, we use the discriminator $f_{\bm{\phi}}$ whose architecture is composed of 1D convolution, residual block, and Leaky ReLU. We set $\lambda_{\text{adv}}=1.0\times 10^{-3}$ as the weight for adversarial training. All our models are trained for 6K epochs with the AdamW optimizer using the batch size of 128 and a cosine-annealing learning rate of $1.0\times 10^{-4}$ with a warm-up. All other settings, including those for the stage 2 model $\epsilon_{\theta}$ training and text embedding $\tau(\bm{c})$, follow the official MLD implementation.
In case (ii), we use motion VQ-VAE whose codebook size is set to $512\times 512$. We utilize the AdamW optimizer, batch size of 256, and exponential moving constant of 0.99. For adversarial training, we use the discriminator $f_{\bm{\phi}}$ whose architecture is composed of 1D convolution, residual block, and Leaky ReLU, the same as in case (i). We train the first 200K iterations with a learning rate of $2.0\times 10^{-4}$, and 100K with a learning rate of $1.0\times 10^{-5}$. All other settings follow the official T2M-GPT implementation.
Dataset Stage 1 model rFID $\downarrow$
HumanML3D VQ-VAE 0.065
VQ-VAE + GAN 0.057
VQ-VAE + SAN 0.051
KIT-ML VQ-VAE 0.665
VQ-VAE + GAN 0.377
VQ-VAE + SAN 0.299
Table 3. Analysis of motion reconstruction performance (training stage 1 with VQ-VAE) on HumanML3D and KIT-ML
@@ 4.2. Implementation Details
MMM (Pinyoanuntapong et al., 2024) $0.504^{\pm.003}$ $0.696^{\pm.003}$ $0.794^{\pm.002}$ $0.080^{\pm.003}$ $2.998^{\pm.007}$ $9.411^{\pm.058}$ $1.164^{\pm.041}$
Continuous (raw data) MotionDiffuse (Zhang et al., 2024) $0.491^{\pm.001}$ $0.681^{\pm.001}$ $0.782^{\pm.001}$ $0.630^{\pm.001}$ $3.113^{\pm.001}$ $9.410^{\pm.049}$ $1.553^{\pm.042}$
@@
Continuous (latent) MLD (Chen et al., 2023) $0.481^{\pm.003}$ $0.673^{\pm.003}$ $0.772^{\pm.002}$ $0.473^{\pm.013}$ $3.196^{\pm.010}$ $9.724^{\pm.082}$ $2.413^{\pm.079}$
MoLA (ours) $0.514^{\pm.003}$ $0.705^{\pm.002}$ $0.797^{\pm.002}$ $0.362^{\pm.009}$ $3.162^{\pm.008}$ $9.672^{\pm.079}$ $2.355^{\pm.121}$
Table 4. Comparison with state-of-the-art methods on HumanML3D.
4.3. Effectiveness of Adversarial Training
We first discuss the effect of adversarial training on Stage 1. Here, we adopt VAE and VQ-VAE described in Section 4.2 as the baselines. The case of VAE is intended to show the effects in a continuous latent space, while the case of VQ-VAE is intended to show the effects in a discrete latent space. The results for VAE combined with GAN/SAN are shown in Table 2, and those of VQ-VAE in Table 3. We utilize the reconstruction FID (rFID) as an evaluation metric. As shown in the tables, the adversarial training is effective in both VAE and VQ-VAE in the motion reconstruction task. In particular, we can improve the performance of the Stage 1 model by adopting the SAN framework instead of the conventional GAN. A better rFID is directly related to the upper bound of the overall performance of a text-to-motion model. Thus, we use VAE trained with the SAN framework for our MoLA. We demonstrate that adversarial training in Stage 1 actually improves the generation performance of the entire model in the next section.
[…]

### v3

@@ 3.2.1. Learning a motion latent representation with VAE-GAN
where $\lambda_{\text{reg}}$ is a hyperparameter for balancing the two terms. The encoder $q_{\bm{\eta}}$ can be trained to produce a low-dimensional latent representation, and the decoder $g_{\bm{\psi}}$ can also be trained to accurately reconstruct the input motion sequences from the latent representations. Besides, we separate the decoder output into two components, denoted as $g_{\bm{\psi}}(\bm{z})=[\bm{m}^{\prime}(\bm{z}),\bm{a}^{\prime}(\bm{z})]$, where $\bm{m}^{\prime}(\bm{z})\in\mathbb{R}^{(N-1)\times L}$ and $\bm{a}^{\prime}(\bm{z})\in\mathbb{R}^{L}$ correspond to the original motion and proposed activation variable. We adopt a binary cross entropy (BCE) loss for the activation variable. The following $\mathcal{J}_{\text{MotionVAE}}$ is used as the loss function for the VAE part:
$\mathcal{J}_{\text{MotionVAE}}(\bm{\psi},\bm{\eta};\bm{x})=~{}\mathbb{E}_{q_{\bm{\eta}}(\bm{z}|\bm{x})}[\|\bm{m}-\bm{m}^{\prime}(\bm{z})\|_{2}^{2}$
(4) $\qquad+\lambda_{\text{act}}\mathcal{L}_{BCE}(\bm{a},\bm{a}^{\prime}(\bm{z})))]+\lambda_{\text{reg}}D_{KL}(q_{\bm{\eta}}(\bm{z}|\bm{x})\|p(\bm{z})),$
where $\lambda_{\text{act}}$ is a hyperparameter for weighting the BCE loss term.
@@
We formulate the overall loss for the VAE as the sum of the negative ELBO and adversarial loss,
(6) $\min_{\phi,\psi}\mathbb{E}_{p(x)}\left[\mathcal{J}_{\text{MotionVAE}}(\bm{\psi},\bm{\eta};\bm{x})+\lambda_{\text{adv}}\mathcal{J}_{\text{GAN}}(\bm{\psi},\bm{\eta};\bm{\phi},\bm{x})\right],$
@@
We train both the VAE and the discriminator with Equations (5) and Equation (6) in an alternating way.
@@ 3.2.2. From GAN- to SAN-based discriminator
We apply the slicing adversarial network (SAN) framework (Takida et al., 2024) to further enhance the motion VAE, based on a prior report showing SAN-based models perform better than the GAN counterparts. The impact of this replacement on motion generation is discussed in Section 4, while a detailed explanation of SAN is provided in the Appendix.
3.2.3. Architectures
We use a standard CNN-based architecture in the motion VAE encoder $q_{\bm{\eta}}$, decoder $g_{\bm{\psi}}$ and discriminator $f_{\bm{\psi}}$, consisting of 1D convolution, a residual block, and Leaky ReLU. For temporal downsampling and upsampling, we use stride 2 convolution and nearest interpolation, respectively. Specifically, the motion sequence $\bm{x}\in\mathbb{R}^{N\times L}$ is encoded into a latent vector $\bm{z}\in\mathbb{R}^{d_{z}\times d_{l}}$ with downsampling ratio of $d_{l}=L/4$. This architecture is inspired by (Esser et al., 2021; Zhang et al., 2023b).
@@ 3.3. Stage 2: Motion Latent Diffusion
3.3.1. Text-conditional motion generation
In this section, we train a text-conditioned diffusion model on the low-dimensional motion latent space obtained by the autoencoder learned in the stage 1 (Section 3.2). Using the trained model, we perform motion generation conditioned on text. First, we define a time-dependent sequence $\bm{z}_{0},\bm{z}_{1},\dots,\bm{z}_{t},\dots,\bm{z}_{T}\in\mathbb{R}^{d_{z}\times d_{l}}$ (starting from the VAE encoder output $\bm{z}_{0}=\bm{z}\sim q_{\bm{\eta}}(\bm{z}|\bm{x})$), which is derived from the following Markov diffusion process in the latent space:
@@ 3.3.1. Text-conditional motion generation
where $T>0$ and the constant $\alpha_{t}\in(0,1)$ is a pre-defined noise-scheduling parameter that determines the forward process. The forward process allows for the sampling of $\bm{z}_{t}$ at an arbitrary time step $t$ in a closed form: $\bm{z}_{t}=\sqrt{\bar{\alpha}_{t}}\bm{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$, where $\bar{\alpha}_{t}\coloneq\prod_{s=1}^{t}\alpha_{s}$ and $\epsilon\sim\mathcal{N}(\bm{0},\bm{I})$.
As our goal is text-to-motion generation, our interest is in the conditional distribution $p(\bm{z}|\bm{c})$ given the text prompt $\bm{c}$. Here, similar to many text-conditioned latent diffusion models (Rombach et al., 2022; Chen et al., 2023), we train the conditional model $\epsilon_{\theta}(\bm{z}_{t},t,\tau(\bm{c}))$ conditioned on the output of a text encoder $\tau(\bm{c})$, using the following objective function:
@@
where $\bm{z}_{0}$ and $\bm{c}$ are drawn from the joint empirical distribution. In addition, as done in prior works, we adopt classifier-free guidance (Ho and Salimans, 2022) and train the model unconditionally, i.e., without a text prompt, with a certain probability during training.
During inference (text-to-motion generation), the trained diffusion model $\epsilon_{\theta}(\bm{z}_{t},t,\tau(\bm{c}))$ is used to generate $\bm{z}_{0}$ through a denoising process conditioned on the text prompt $\bm{c}$. We adopt the sampling scheme of DDIM (Song et al., 2021a) with trailing sample steps (Lin et al., 2024), in which each sampling step is defined as:
@@
In actual inference, we use the estimated $z_{0}$ and the VAE decoder trained in stage 1 to obtain $g_{\bm{\psi}}(\bm{z}_{0})$ as the generated motion sequences. Variable length motion generation is then achieved by clipping a part of $g_{\bm{\psi}}(\bm{z}_{0})$ where the activation variable $a^{i}$ introduced in Section 3.1 satisfies $a^{i}<\delta\ (0<\delta<1)$. The effect of this approach is discussed in Section 4.1.
3.3.2. Architecture
We employ a diffusion transformer (DiT)-based architecture in our stage 2 model. The transformer used follows a standard structure of stacked blocks consisting of an attention layer and gated multilayer perceptrons (MLP) connected in series, with skip connections around each. Additionally, layer normalization is employed on the inputs of both the attention layer and the MLP. At the input and output of the transformer, linear mapping is used to convert from the latent dimension of the stage 1 model to the embedded dimension of the transformer. Text CLIP embedding is also added as input to the transformer, along with an embedding describing the current time step of the diffusion process. This architecture is inspired by (Evans et al., 2024), which established SOTA performance in audio generation.
@@ 3.4. Controllable Motion Generation on Latent Diffusion Sampling
which is derived from Bayes’ rule. The conditional generation based on the aforementioned property can be regarded as a sequential procedure implemented as follows. First, a denoised sample $\bm{z}_{t-1}$ is obtained from $\bm{z}_{t}$ by the sampling step in Equation (10), without considering the given condition $\bm{y}$ (the first term in Equation (10)). Subsequently, $\bm{z}_{t-1}$ is further updated using the gradient of the log-likelihood term with respect to $\bm{z}_{t}$ (the second term).
The challenge here is that the log-likelihood term is computed based on noisy samples $\bm{z}_{t}$. In the classifier-guided diffusion (Song et al., 2021b), time-dependent classifiers for $\bm{z}_{t}$ have to be trained, which requires additional training. In contrast, in training-free methods such as DPS, FreeDoM, and MPGD, this term is approximated with the current clean data estimate $\bm{z}_{0|t}$ and a loss function $\mathfrak{L}(\bm{x};\bm{y})$ defined for clean data. This loss function can be flexibly set depending on the task. For example, in inverse problems of the form $\bm{y}=\mathcal{A}(\bm{x})$, where $\mathcal{A}$ is a differentiable function with respect to $\bm{x}$, the loss function can be set as $\|\bm{y}-\mathcal{A}(\bm{x}_{0|t})\|_{2}^{2}$, where $\bm{x}_{0|t}=g_{\bm{\psi}}(\bm{z}_{0|t})$ is a clean data estimate in the original data domain and obtained through the VAE decoder. In FreeDoM’s FaceID-guided generation, the $\ell_{2}$ distance of features obtained by inputting facial images into a facial recognition model is used as the loss.
Here, we adopt MPGD (He et al., 2024), a fast yet high-quality guidance method applicable to latent diffusion models. Following the denoising step, it updates the denoised sample based on the loss function $\mathfrak{L}(\bm{x};\bm{y})$ as follows:
(12) $\bm{z}_{t-1}\leftarrow\bm{z}_{t-1}-\rho_{t}\sqrt{\bar{\alpha}_{t-1}}\nabla_{\bm{z}_{0|t}}\mathfrak{L}(g_{\bm{\psi}}(\bm{z}_{0|t});\bm{y}),$
@@
More specifically, in editing motion tasks we generate motions to match given specific poses or trajectory control signals. To deal with various motion editing tasks in this framework, the loss function is designed as follows:
(13) $\mathfrak{L}_{\text{Motion}}(g_{\bm{\psi}}(\bm{z}_{0|t});\bm{y})=\Sigma_{n}\Sigma_{l}m_{nl}\|\mathfrak{R}(g_{\bm{\psi}}(\bm{z}_{0|t}))_{nl}-y_{nl}\|_{2},$
@@
where $n$ and $l$ are indices of joint and frame, respectively, and $m_{nl}$ is a binary value indicating whether the control position $y_{nl}$ contains a valid value at frame $l$ for joint $n$, and $\mathfrak{R}(\cdot)$ is a function that converts the motion features including the joint’s local positions to global absolute locations. We set $\mathfrak{L}_{\text{Motion}}$ for loss function $\mathfrak{L}$ in the guided generation to measure the distance between desired constraints $\bm{y}$ and the joint locations of the generated motion. Target locations as constraint $\bm{y}$ can be specified for any subset of joints in any subset of motion frames. [^2] As examples of motion editing, if $\bm{y}$ is given as the start-end positions, we can handle the motion in-betweening. If $\bm{y}$ is given as the lower body positions, we can edit the upper body corresponding to the lower one. If $\bm{y}$ is given as the pelvis trajectory, it corresponds to the path-following task. We leave the task details to Section 4.2. Note that the guided generation framework in Equations (12) and Equation (13) has the potential to generate motion while dealing with a variety of time and spatial constraints not limited to these three task examples. Editing a generated motion to match specific poses or follow a specific trajectory is achieved by minimizing $\mathfrak{L}_{\text{Motion}}$ using the update rule in Equation (12).
@@
MMM (Pinyoanuntapong et al., 2024b) $0.504^{\pm.003}$ $0.696^{\pm.003}$ $0.794^{\pm.002}$ $0.080^{\pm.003}$ $2.998^{\pm.007}$ $9.411^{\pm.058}$ $1.164^{\pm.041}$
ParCo (Zou et al., 2024) $0.515^{\pm.003}$ $0.706^{\pm.003}$ $0.801^{\pm.002}$ $0.109^{\pm.003}$ $2.927^{\pm.008}$ $9.576^{\pm.088}$ $1.382^{\pm.060}$
BAMM (Pinyoanuntapong et al., 2024a) $0.525^{\pm.002}$ $0.720^{\pm.003}$ $0.814^{\pm.003}$ $0.055^{\pm.002}$ $2.919^{\pm.008}$ $9.717^{\pm.089}$ $1.687^{\pm.051}$
Continuous (data space) MotionDiffuse (Zhang et al., 2024) $0.491^{\pm.001}$ $0.681^{\pm.001}$ $0.782^{\pm.001}$ $0.630^{\pm.001}$ $3.113^{\pm.001}$ $9.410^{\pm.049}$ $1.553^{\pm.042}$
@@
Continuous (latent) MLD (Chen et al., 2023) $0.481^{\pm.003}$ $0.673^{\pm.003}$ $0.772^{\pm.002}$ $0.473^{\pm.013}$ $3.196^{\pm.010}$ $9.724^{\pm.082}$ $\bm{2.413^{\pm.079}}$
MotionLCM (Dai et al., 2024) $0.502^{\pm.003}$ $0.698^{\pm.002}$ $0.798^{\pm.002}$ $0.304^{\pm.012}$ $3.012^{\pm.007}$ $9.607^{\pm.066}$ $2.259^{\pm.092}$
MoLA (ours) $\bm{0.516^{\pm.006}}$ $\bm{0.712^{\pm.005}}$ $\bm{0.805^{\pm.004}}$ $\bm{0.115^{\pm.004}}$ $\bm{3.008^{\pm.016}}$ $9.885^{\pm.152}$ $2.156^{\pm.157}$
Table 1. Comparison with state-of-the-art methods on HumanML3D (Guo et al., 2022). Note that discrete representations do not allow for training-free motion editing; therefore, methods based on VQ-based latent representations (Discrete) are grayed out. The best scores for each metric in the methods using VAE-based latent representations (Continuous (latent)) are highlighted in bold.
4. Experiments
We evaluate the performance of MoLA on two tasks: motion generation (Section 4.1) and motion editing (Section 4.2). Our results demonstrate that MoLA achieves its three key objectives: (1) fast and high-quality generation, (2) variable-length generation, and (3) multiple motion editing tasks in a training-free manner. To validate these objectives, we utilize the HumanML3D (Guo et al., 2022).333 This dataset contains 14,616 human motions from the AMASS (Mahmood et al., 2019) and HumanAct12 (Guo et al., 2020) datasets and 44,970 text descriptions. Details of the experimental setup as well as additional results on KIT-ML (Plappert et al., 2016) are provided in the Appendix.